40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog, Page 3 of 54  Not logged in ELOG logo
ID Date Author Type Category Subject
  2596   Tue Jun 29 17:16:29 2021 Ian MacMillanDailyProgressCryo vacuum chamberTemp Controller

Radhika and I were able to connect to the temperature controller using telnet. We simply reassigned the IP address of the temp controller to fit in the IP address range of the router/network. i.e. if the router is 132.232.114.1 then the temp controller might be 132.232.114.112. This shows that the router and network are working correctly again.

  2595   Tue Jun 29 15:45:43 2021 Ian MacMillanDailyProgressGeneralNetwork Issue

I was able to connect to the router (the one on top of the switch in the rack next to the orange tool drawers) and reconfigure it using the same internal IP as found labeled and mentioned in [2201]. The router now has a different external IP that I wrote on the router (twice). The network/router and switch is now working in the lab.

Update: It turns out that the router reset itself somehow. so any settings/ port forwarding that was set up before has probably been wiped out and would need to be set up again.

SSHing into the QIL needs to be set up again. 

  2594   Fri Jun 25 09:47:33 2021 RadhikaDailyProgressCryo vacuum chamberCooldown data

I've attached cooldown data from 6/21-6/24. Unexplicably, the CTC-100 stopped logging temperatures at ~50 hours after the start of cooling, or around 5pm Wednesday. The red logging indicator was definitely on at 3pm Wednesday, but it was off when I checked in Thursday afternoon. It must have somehow been disabled in between. On the bright side, the temperature of the workpiece stabilized at 66K. I had checked in at 53 hrs, 62 hrs, and 64 hrs and noted the temperatures by eye (green data points on attachment 2). 

The time constant for the workpiece is ~ 36.4 hrs, about the same as the previous run (tau = -40/ln(100/300)). While no improvements were made on this front (cooldown still very slow), it stabilized 6K below the previous run. 

The RTD at the cold head recorded a shoot-up in temperature right at the beginning of cooldown, which makes Stephen/me think that it debonded from the cold head. This temp reading would have been informative for our model, but we will verify it's status when we open up. We can find a better bonding mechanism for the next cooldown.

[Update] I let the chamber warm up over the weekend, and brought it up to room pressure on Monday 6/28. Next steps are to open up and investigate the cold head RTD and make improvements to bonding mechanism, if necessary. We will assess any other improvements (model driven) before the next cooldown.

Attachment 1: cooldown_6-21.pdf
cooldown_6-21.pdf
Attachment 2: cooldown_model_vs_data_v2.pdf
cooldown_model_vs_data_v2.pdf
  2593   Wed Jun 23 16:25:49 2021 Ian MacMillanDailyProgressGeneralNetwork Issue

I accidentally powered down and restarted the TP-link safe stream router and now nothing will connect to it. Not even things that are plugged into it. I can't even connect to the router's setup screen. I have no idea how to solve this issue short of resetting the router. I don't know how it was configured before so I will not reset it because I don't know how it was configured.

  2592   Wed Jun 23 15:03:12 2021 Ian MacMillanDailyProgressCDSTesting DAC2ADC code

I brought DAC2ADC_test.py code from the 40m to test in the QIL. I added a parser arg for the matrix root name (see attached code). I am running into the same problem as I did at the 40m where the channels seem to be locked to their values. I attached (attachment 2) the results text file that shows that for all inputs the outputs are the same. In the GDS screen the ADC is in red but I'm not sure how it got there or how to fix it.

run the test with these commands in the QIL

 $  cd IanMacMillan
 $  python DAC2ADC_Test.py -e 'C4:TST' -m 'cdsMuxMatrix'

One improvement that we could make to the code is have it zero out all of the matrix elements before it starts.

Attachment 1: QIL_CDS_TEST.tar.gz
  2591   Mon Jun 21 10:59:41 2021 RadhikaDailyProgressCryo vacuum chamberChamber updates 6/16-6/18

[Radhika, Stephen]

Background: There are currently 4 available RTDs for the large cryo chamber (henceforth named Megastat). We originally had 1 for the workpiece, 1 for the heater, 1 for the inner shield, and 1 for the outer shield. During the last sessions with the chamber open, the RTDs were moved around. Now there is 1 for the workpiece, 1 for the baseplate (bottom lid), one for the outer shield, and 1 for the cold head. These locations are marked in the diagram (attachment 9). 

On Wednesday 6/16, we opened up the chamber and removed the shields and coldplate. We added foil to the inner surface of the bottom lid (attachment 10) and attached an RTD there (we forgot to take a picture). Stephen made a new foil collar, and we decided to push it against the chamber walls as a better alternative to having the foil touch the coldplate (attachment 6). We added an extra layer of aluminized mylar to wrap the copper braid, and we fed another RTD through the braid tube to be attached to the coldhead. We lastly placed the coldplate and shields back in place.

On Friday 6/18, I attached the remaining RTDs (coldhead, outer shield, workpiece). We decided to remove the aluminum foil previously covering the coldhead, due to of fear of shorting to the tube wall. I taped several pieces of mylar together to cover the coldhead and insulate it from the wall (attachments 1-3). I placed the mylar contraption into the T (attachment 4) and then closed the bottom flange. I placed back the shield lids and the main lid of the chamber (attachments 7-8). I tried to pump down, but the pressure was stabilizing on the order of 1e-1 torr. Unsure of why the pressure wasn't decreasing, I turned off the pump and left for the day.

Today 6/21, I touched base with Stephen and we realized I forgot to replace the copper gasket in the bottom flange of the T. I then unscrewed and replaced the gasket. I pumped down and the pressure reached on the order of 1e-4 torr, so I proceeded with cooling. After a few hours I could see that the cold head RTD was reading a temperature around 180K. We will extract and plot cooling data late tomorrow or Wednesday.

Lastly, a chamber diagram is attached (attachment 9). The 4 RTD locations are marked by a red 'x'. The chamber components are numbered in blue (detailed below):

1. Workpiece

2. Heater

3. Inner shield

4. Outer shield

5. Copper braid (wrapped with mylar)

6. Cold head

7. Coldplate

8. Baseplate (bottom lid)

Attachment 1: IMG_1735.jpeg
IMG_1735.jpeg
Attachment 2: IMG_1738.jpeg
IMG_1738.jpeg
Attachment 3: IMG_1737.jpeg
IMG_1737.jpeg
Attachment 4: Photo_on_6-18-21_at_15.36.jpeg
Photo_on_6-18-21_at_15.36.jpeg
Attachment 5: Photo_on_6-18-21_at_16.32.jpeg
Photo_on_6-18-21_at_16.32.jpeg
Attachment 6: IMG_8947.jpeg
IMG_8947.jpeg
Attachment 7: Photo_on_6-18-21_at_16.33.jpeg
Photo_on_6-18-21_at_16.33.jpeg
Attachment 8: Photo_on_6-18-21_at_16.33_#2.jpeg
Photo_on_6-18-21_at_16.33_#2.jpeg
Attachment 9: New_Note.png
New_Note.png
Attachment 10: IMG_8973.jpeg
IMG_8973.jpeg
  2590   Fri Jun 18 10:15:14 2021 StephenUpdateCryo vacuum chamberTemperature sensor considerations

RTD thoughts - we have just been using the sensors that were provided, without noticing their constraints or deficiencies.

  • our RTD p/n is 615-1123-ND (Digikey), which is a Littlefuse PPG102A6 platinum RTD and has an apparent temperature range of -200 °C (73 K) to + 600 °C. The primary data sheet does not have a Resistance vs Temperature curve, simply presenting the slope of the temperature dependence in a parameter "Temperature Coefficient of Resistance" (TCR), but a table is available in an auxiliary document called "RT Chart". An image of this chart is attached below.
  • we have not violated that 73 K lower limit yet, but we are about to as we mount an RTD directly to the coldhead. Let's see how that RTD on the coldhead will withstand the lower temperatures.
  • the CTC100 manual indicates that any arbitrary calibration curve may be input, but there are a number of sensors' calibration curves built in - it might be a good idea to make our next sensor decision based on that list.
  • there is a Lakeshore guide and category page for temperature sensor selection - reconfiguring our 4 RTDs would cost over a kilobuck through Lakeshore, but perhaps we can learn general ideas from the guide as well.
    • it seems that switching to their Cernox line would be helpful in terms of packaging options, and would be the most accurate.
    • their line of silicon diodes would be suitable and has flexible packaging as well.
    • their platinum RTDs also have low temperature range and would be suitable. Packaging is cylindrical, so might be best to pursue the aluminum housing with a bolt hole.

Planning for next steps:

  • for now, it seems that we could get by with our generic and 73 K limited RTDs, and this option is tempting as it requires no additional effort.
  • if we decide we really want to have reliable sensing down to sub-50 K temperatures, we should move to one of the Lakeshore product lines (hopefully one which the CTC100 is configured with a calibration curve for) for about a kilobuck.
  • we should engage in more serious sensor design before Mariner, regardless of whether we take any action now.
    • as a starting point, the Lakeshore catalog and appendices (ref. product info page) and other resources should be absorbed, for considerations like thermal anchoring, lead length, benefits of 4-lead wiring, polyimide leads causing less conduction to the sensor than teflon, phosphor bronze having lower thermal conduction than copper, etc. Most of these topics are gathered from Appendix C.
Attachment 1: RTD_Resistance_vs_Temperature_for_PPG102A6.png
RTD_Resistance_vs_Temperature_for_PPG102A6.png
  2589   Wed Jun 16 17:17:12 2021 KojiUpdateGeneralI2 cell

I was searching an I2 (Iodine) cells back to the days of the laser gyro.

I found a likely box at a very tricky location. Took the photos and returned to this tricky place.
 

2021/Jul The box was moved to the OMC lab (KA)

Attachment 1: P_20210616_170104.jpeg
P_20210616_170104.jpeg
Attachment 2: P_20210616_170038.jpeg
P_20210616_170038.jpeg
Attachment 3: P_20210616_170021.jpeg
P_20210616_170021.jpeg
  2588   Fri Jun 11 16:48:31 2021 Aidan, StephenSummary2um PhotodiodesChamber is leaking

We hit another dead-end with leak hunting the IR labs dewer (we replaced screws and helicoil on the valve connection but there is still a big leak). We cleaned the flange and O-ring with isopropanal and replaced the threads with helicoil but still get the same sort of leak where we only hit 1E-2 Torr after 5 minutes of pumping and stablize around 1E-3.

After turning off the pumping station, the pressure rose quickly to 1Torr (in roughly 10 minutes or so).

 

Quote:

[Stephen, Aidan, Wednesday 09 June]

Summary and Plan:

  • Poor sealing at the output valve (The Gap) needs to be resolved.
  • Planning to install #4-40 helicoils today (chamber will remain sealed, will need to remove output valve and cover output orifice, then transport the chamber to the WB EE shop for redrilling of holes.)
  • Meeting with Nina and Aidan this afternoon to iterate one more time.

Troubleshooting steps taken:

  1. Aidan took us through the full sequence of pump down and disassembly to bring me up to speed.
  2. We opened the lid and inspected the old o-ring.
    1. Signs of plastic deformation and of small flecks of particulate near sealing surface - good idea to change.
  3. We found new o-rings in a box from the Cryo lab, and one of these was swapped in after a good wipedown with IPA.
  4. Upon pumpdown, Aidan compared behavior and found no meaningful change to rate of pumpdown or stable pressure in e-4 torr range after 10+ minutes.
    1. By valving off [chamber + gauge] from pump line, it was clear that there was a leak in that volume, as within seconds the pressure rose from e-4 torr to e-2 torr, and stabilized at _(need to confirm - e0?)_ torr over ~10 minutes.
  5. Attempted to squirt IPA along o-ring seals, but there was not good access to the sealing surfaces, so this was a null test
  6. Looked closer at all of the chamber features, and noticed The Gap between chamber wall and chamber output valve, pictured in [Attachment 1]. Not good! But promising as a leak source.
    1. Three of the four screws were found to be loose due to apparent thread damage.
  7. IPA was squirted into The Gap at stable pressure of e-4 torr, but no change in pressure was noticed.
  8. Longer #4-40 screw reinstallation was attempted, and I could feel a small amount of pull at the very tip of the screw, but tightening the screws led to that small pull to fail as well - need to rework.
  9. The Clamp was installed and The Gap was closed [Attachment 2].
    1. When isolated from the vacuum pump, the chamber pressure progressed more slowly. Within seconds, we were at e-3 torr, and over 10 minutes the pressure stabled at e-2 torr, about 30x lower pressure per Aidan's records.

 

 
Attachment 1: IMG_3002.jpg
IMG_3002.jpg
Attachment 2: IMG_3001.jpg
IMG_3001.jpg
  2587   Fri Jun 11 16:00:31 2021 RadhikaDailyProgressCryo vacuum chamberLarge chamber heating

Summarizing heater trends from Wednesday, 6/9. Reiterating from post [2584], I ran the CTC-100 temp controller's auto tune routine to adjust PID coefficients for the heater. I then set the setpoint to 123K (operating temperature); the controller takes in the workpiece temperature as feedback. The heating data for this period is attached.

It took under 30 minutes for the temperature to rise from 72K (current cooling limit) to within a degree of the setpoint, 123K. The power delivered to the heater (bottom plot of first attachment) stayed below 25W, the limit we hardcoded. The workpiece temperature rises smoothly and plateaus around the setpoint, without significant overshooting. The controller holds the setpoint temperature pretty well thereafter. The second attachment is zoomed in on the workpiece temperature alone.

This run served as a test of the temperature controller's stability at the desired setpoint. Moving forward, we will continue to improve the cooling capacity of the chamber guided by our model. Once we optimize how cold we can get, we now know the heater can hold us at the desired temperature setpoint.

Attachment 1: heating_6-9.pdf
heating_6-9.pdf
Attachment 2: heating_6-9_workpiece.pdf
heating_6-9_workpiece.pdf
  2586   Fri Jun 11 07:42:54 2021 StephenSummary2um PhotodiodesChamber is leaking

[Stephen, Aidan, Wednesday 09 June]

Summary and Plan:

  • Poor sealing at the output valve (The Gap) needs to be resolved.
  • Planning to install #4-40 helicoils today (chamber will remain sealed, will need to remove output valve and cover output orifice, then transport the chamber to the WB EE shop for redrilling of holes.)
  • Meeting with Nina and Aidan this afternoon to iterate one more time.

Troubleshooting steps taken:

  1. Aidan took us through the full sequence of pump down and disassembly to bring me up to speed.
  2. We opened the lid and inspected the old o-ring.
    1. Signs of plastic deformation and of small flecks of particulate near sealing surface - good idea to change.
  3. We found new o-rings in a box from the Cryo lab, and one of these was swapped in after a good wipedown with IPA.
  4. Upon pumpdown, Aidan compared behavior and found no meaningful change to rate of pumpdown or stable pressure in e-4 torr range after 10+ minutes.
    1. By valving off [chamber + gauge] from pump line, it was clear that there was a leak in that volume, as within seconds the pressure rose from e-4 torr to e-2 torr, and stabilized at _(need to confirm - e0?)_ torr over ~10 minutes.
  5. Attempted to squirt IPA along o-ring seals, but there was not good access to the sealing surfaces, so this was a null test
  6. Looked closer at all of the chamber features, and noticed The Gap between chamber wall and chamber output valve, pictured in [Attachment 1]. Not good! But promising as a leak source.
    1. Three of the four screws were found to be loose due to apparent thread damage.
  7. IPA was squirted into The Gap at stable pressure of e-4 torr, but no change in pressure was noticed.
  8. Longer #4-40 screw reinstallation was attempted, and I could feel a small amount of pull at the very tip of the screw, but tightening the screws led to that small pull to fail as well - need to rework.
  9. The Clamp was installed and The Gap was closed [Attachment 2].
    1. When isolated from the vacuum pump, the chamber pressure progressed more slowly. Within seconds, we were at e-3 torr, and over 10 minutes the pressure stabled at e-2 torr, about 30x lower pressure per Aidan's records.
Attachment 1: IMG_8865.JPG
IMG_8865.JPG
Attachment 2: IMG_8881.JPG
IMG_8881.JPG
  2585   Thu Jun 10 15:13:26 2021 RadhikaDailyProgressCryo vacuum chamberCooldown modeling vs. data

Code for simple heat transfer modeling can be found here: https://git.ligo.org/voyager/mariner40/-/blob/master/CryoEngineering/qil_simple_heat_transfer.ipynb

My original 1D cooldown script modeled conductive cooling along the copper braid as: Pcool = k * A/L * (T - T_set); where Pcool is the cooling power, k is the thermal conductivity of copper, A is the cross-sectional area of the braid, and L is the length of the braid. I changed the code to instead use the tabulated power vs. temperature points for the CH-104 coldhead, taken from Paco's script qil_heat_estimate.ipynb. The first figure compares the interpolated curves from the tabulated values (at 50Hz and 60Hz operation), to the original conductive transfer model (50K setpoint). The original conductive power-temp relationship is linear, which overestimates the cooling power at high temperatures. Switching to the tabulated points results in more realistic model. Moving forward, I intend to use the 50Hz interpolated curve. 

The script considers radiative heating to the coldplate from the the chamber bottom (rough aluminum) and the outer shield (coated in aluminum foil). It assumes over a long period of time that the inner shield and coldplate temperatures are equal.

The second figure shows the results of this model alongside the actual coolddown data extracted from the CTC-100. It is clear that the model is not accounting for additional radiative heat sources that would explain the slower cooldown and higher final temperature. Adding in model complexity is my current focus. 

 

Attachment 1: coldhead_capacity.pdf
coldhead_capacity.pdf
Attachment 2: cooldown_model_vs_data.pdf
cooldown_model_vs_data.pdf
  2584   Wed Jun 9 13:35:07 2021 RadhikaDailyProgressCryo vacuum chamberHeater actuation

Today I attempted to auto tune PID coefficients for the heater, so that we can reach and maintain a setpoint of 123K with appropriate ramp-up. The workpiece was around 72K originally. For auto tuning, I set the max power for the heater to 25W. I adjusted the lag time to 30s, and changed the setpoint to 72K so that the tuning response measured how stable the system is when being perturbed from our set point. The auto tune process ran without an error; however, by default the tuning mode switched to "step" tuning, and I am not sure why this occurred. The tuning took roughly 5 minutes to complete. The final message is attached; the adjusted parameters were:

Gain: 2.703

Lag: 125.8 s

Time constant: 271.9 s

I was expecting the adjusted output to be new PID coefficients, but I noticed that the PID coefficients did seem to change after this process. To begin warmup to 123K, I changed the setpoint to 123K and let the heater do its thing (the feedback temperature is set to that of the workpiece). The heater power stabilized to around 17W, and the temperature of the workpiece reached within a degree K of the setpoint within 30 minutes. I am letting the temperature hold at 123K overnight and plan to return tomorrow morning to check on it and extract the heating data for plotting. 

Attachment 1: IMG_1722.jpeg
IMG_1722.jpeg
  2583   Tue Jun 8 12:58:40 2021 StephenDailyProgressCryo vacuum chamberMaximum power for temperature actuation

Quick log establishing the maximum power for our thermal actuation:

Heater: HSA25100RJ from TE, unknown sourcing. Acetone wiping cleaned off p/n and markings from body, should engrave at next opportunity, but [Attachment 1] from many months ago shows the p/n. Note that this is not the current mounting configuration - [Attachment 2] is more similar to current mounting. Anyway, according to the datasheet (now added to the QIL wiki at Documentation > Manuals) this heater is rated for 25W and has a resistance of 100Ω.

Leads: unknown, and not super important unless we had tiny hair conductor - I am not in lab presently, but it appears from our connector (Lesker FTACIR19AC) that we must have 20-24 AWG,

Carrying the 7, the current through the Heater will be 0.25 A at max actuation, and the 20-24 AWG insulated copper leads will have plenty of ampacity for this load (plus, they are cooled, so normal current capacity considerations fly out the window a bit).

Conclusion: 25W actuation will be the limit that we will apply in the CTC100 temperature actuation routine.

From first trials yesterday, the response at ~20W (at starting temperatures around 80K) appears to be on the order of 1 degree per minute, which should be just fine for actuating to maintain a +/-  1 degree constant setpoint with static thermal loads. More to follow on trials implementing temperature control.

Note that the QIL Wiki points to the DCC (which contains a budget that was helpful resources to trace these purchases), the datasheets and other documentation, and also points to the QIL Cryo Vacuum Chamber photo album, which hosts the images below.

Attachment 1: IMG_8495.JPG
IMG_8495.JPG
Attachment 2: IMG_8745.JPG
IMG_8745.JPG
  2582   Fri Jun 4 16:15:00 2021 RadhikaDailyProgressCryo vacuum chamberCooldown from 6/2-6/4

I extracted cooldown data from the CTC100 USB around 1pm today (~50 hours of cooldown). I've attached a log plot below. The heater RTD seemed to be behaving weirdly at the beginning, but soon stabilized and cooled as expected.

I estimated the time constant for the workpiece: 50 hr/ (300K - 90K) = 0.24 hr/K =  ~860 s/K

 

 

Attachment 1: cooldown_6-2_6-4.png
cooldown_6-2_6-4.png
  2581   Wed Jun 2 20:06:07 2021 ranaDailyProgressCryo vacuum chamberWeekend cryo chamber cooling

I'll be curious to see the results of Radhika's thermal model - I am suspicious of this thermal strap contact to the base plate. It would be good if we could instead make a copper mating plate:

  1. groovy milled out loop holder for the strap to go into such that the contact surface is larger
  2. flat base to get smushed into the colde plate
  3. thru holes so that it can be directly screwed down with some washers
  4. screwed down with a torque wrench so that we know that its really tight (i.e. "I think its pretty tight" doesn't really work when we want a good thermal contact)
  5. someone with a MechE degree calculates the proper torque such that we plasticly deform the copper and get it to sit more flush to the cold plate
  6. a cap for the copper plate that really smooshes the copper strap into the bottom part of the mating plate.
  2580   Mon May 31 11:38:13 2021 RadhikaDailyProgressCryo vacuum chamberWeekend cryo chamber cooling

[Stephen, Radhika]

After Thursday's work, we resumed on Friday and lifted up the coldplate to access the collar and baseplate. Stephen added metallized mylar wraps to the peek cylindrical spacers [pic 9, 10]. He took out the cylindrical collar and baseplate (previously there to minimize contact between the colplate and chamber bottom) and replaced them with a foil collar [pic 13, note this is before pushing down the foil over the PEEK spacers]. We re-inserted the coldplate.

In order to wrap/insulate the copper braid, we inserted 2 sheets of metallized mylar into the vacuum tube to surround the braid [pic 11, 14]. We cut the mylar wrapping so that it did not short to the outer shield. We also confirmed that the copper braid is not shorting to the inner shield hole (there was clearance for the shield to be lifted before hitting the braid). The mylar extends to the coldhead, which we then wrapped with aluminum foil, making sure not to short to the walls of the tube [pic 12]. 

We switched the foil wrapping from the inside to the outside of the outer shield, so that any radiative transfer from the inner shield to the outer shield would be absorbed as much as possible (not reflected back). We placed both shields back and bolted down the copper braid loop and workpiece. We re-attached the RTDs, then placed in both shield lids (without bolting them down) and stopped for the day.

Today I checked on the chamber setup and closed up. I started the vacuum pump and it made abnormally loud noises, indicating something was off. The percentage sign continued to flash, indicating that the pump was not reaching 80% speed. I performed 2-3 power cycles, which did not solve the issue. We will pick up to debug the issue early this week.

[update 02 June 2021]

Pumpdown and cooldown were successfully started this morning - Radhika retightened the green leak valve and pumps started just fine. We will check trends on Friday and likely will allow cooldown to proceed over the weekend.

We are both working on adding all of our photos to the photo dump at the ligo.wbridge QIL Cryostat Photo Album.

Attachment 1: IMG_1681.jpeg
IMG_1681.jpeg
Attachment 2: IMG_1684.jpeg
IMG_1684.jpeg
Attachment 3: IMG_1686.jpeg
IMG_1686.jpeg
Attachment 4: IMG_1687.jpeg
IMG_1687.jpeg
Attachment 5: IMG_1688.jpeg
IMG_1688.jpeg
Attachment 6: IMG_1690.jpeg
IMG_1690.jpeg
Attachment 7: IMG_1693.jpeg
IMG_1693.jpeg
Attachment 8: IMG_1695.jpeg
IMG_1695.jpeg
Attachment 9: IMG_8796.JPG
IMG_8796.JPG
Attachment 10: IMG_8805.JPG
IMG_8805.JPG
Attachment 11: IMG_8811.JPG
IMG_8811.JPG
Attachment 12: IMG_8815.JPG
IMG_8815.JPG
Attachment 13: IMG_8803.JPG
IMG_8803.JPG
Attachment 14: IMG_8810.JPG
IMG_8810.JPG
  2579   Thu May 27 16:48:35 2021 RadhikaDailyProgressCryo vacuum chamberWeekend cryo chamber cooling --> started warmup

This morning I vented and opened up the chamber to add aluminum foil and other insulation. Stephen's order of mylar sheets, peet sheets, and G10 rings came in today and I picked it up from Downs.

⁃ unscrewed the bottom conflat of the tee and added foil where I could reach (junction of cold head and copper braid) [pic1]

- cut narrow rings of 0.01'' thickness peet sheet and inserted into the main hole of the inner/outer shields, to prevent shorting to the copper braid [pic 6, 2, 3]

- added foil to the inside of the outershield and outside of the inner shield, poking holes for all viewports [pics 4,5]

⁃ unclamped the copper braid from the coldplate and the RTDs from all locations (shields, heater, workpiece mount) to prepare for lining baseplate with foil [pic 7]

Tomorrow we plan to wrap the shaft of the copper braid with metallized mylar and an additional layer of foil. I left both shields outside of the chamber so that tomorrow we are ready to remove the cold plate and add foil below. 

 

Quote:
  • Cooldown complete as of 4:41, with part 2 of data posted to CSVlogs Box folder with file name "cooldown part 2 20210525.CSV" for Radhika's analysis.
  • Resumed USB logging and started Warmup. Simply turned off cryocooler using compressor panel's green power switch, then flipped circuit breaker lever to off position.
    • Expected to be ready for opening, with warmer-than-freezing temperatures throughout, on Thursday morning.
    • Noticed a "burp" of the gauge pressure readings up to e-2 torr range immediately upon turning off cryocooler, lasting about 1 minute before gauge readings were back down to e-6 range. Interesting, and something to keep an eye on.
  • Linked Active Ion Gauge (5e‐2 to 5e‐10 torr per AIG manual) to Active Pirani Gauge (atm to 7e‐3 torr per APG manual) so that the Edwards TIC controller will enable the high-vacuum AIG during pump down and disabled during venting. Prior, the AIG had to be manually enabled at suitable operating pressures. Followed instructions at Section 4.11 of the TIC controller manual. Recall that all manuals are hosted on the QIL Cryo Vacuum Chamber wiki.

Radhika, feel welcome to post full cooling data to this entry, or to your original - up to you!

 

Attachment 1: IMG_1678.jpeg
IMG_1678.jpeg
Attachment 2: Photo_on_5-27-21_at_15.53.jpg
Photo_on_5-27-21_at_15.53.jpg
Attachment 3: Photo_on_5-27-21_at_16.01.jpg
Photo_on_5-27-21_at_16.01.jpg
Attachment 4: Photo_on_5-27-21_at_16.22.jpg
Photo_on_5-27-21_at_16.22.jpg
Attachment 5: Photo_on_5-27-21_at_16.35.jpg
Photo_on_5-27-21_at_16.35.jpg
Attachment 6: Photo_on_5-27-21_at_16.38.jpg
Photo_on_5-27-21_at_16.38.jpg
Attachment 7: Photo_on_5-27-21_at_16.37.jpg
Photo_on_5-27-21_at_16.37.jpg
  2578   Wed May 26 18:38:25 2021 AidanSummary2um PhotodiodesChamber is leaking

I tried Krytox around the O-ring and also tightening the screws around the valve. The leaking persists at roughly the same rate.

Quote:

I tried pumping down the JPL PD chamber to test the new PD at cryo temperatures. Unfortunately, the chamber can;t get past about 6E-3 Torr with the pump on. As soon as I turned off the pump the pressure rose to around 2 Torr over 20 minutes or so.

I extricated the chamber from the pedestals, flipped it and removed the bottom plate. I cleaned the O-ring with isopropanol and wiped down the mating surface on the chamber (also with iso). I replaced the plate and tightened the screws. Then I returned the chamber to the table and reconnected it to the vacuum system. I tried pumping down once again but I saw pretty much exactly the same situation as before (pressure bottoming out around 6E-3 Torr and then rising quickly again when the pump was turned off).

I guess it's possible that the O-ring is damaged - although I couldn't see anything obivous. We didn't mess around with the viewport (when we replaced the diode a few weeks ago) so I'm hoping there is no issue there.

 

  2577   Tue May 25 16:38:17 2021 StephenDailyProgressCryo vacuum chamberWeekend cryo chamber cooling --> started warmup
  • Cooldown complete as of 4:41, with part 2 of data posted to CSVlogs Box folder with file name "cooldown part 2 20210525.CSV" for Radhika's analysis.
  • Resumed USB logging and started Warmup. Simply turned off cryocooler using compressor panel's green power switch, then flipped circuit breaker lever to off position.
    • Expected to be ready for opening, with warmer-than-freezing temperatures throughout, on Thursday morning.
    • Noticed a "burp" of the gauge pressure readings up to e-2 torr range immediately upon turning off cryocooler, lasting about 1 minute before gauge readings were back down to e-6 range. Interesting, and something to keep an eye on.
  • Linked Active Ion Gauge (5e‐2 to 5e‐10 torr per AIG manual) to Active Pirani Gauge (atm to 7e‐3 torr per APG manual) so that the Edwards TIC controller will enable the high-vacuum AIG during pump down and disabled during venting. Prior, the AIG had to be manually enabled at suitable operating pressures. Followed instructions at Section 4.11 of the TIC controller manual. Recall that all manuals are hosted on the QIL Cryo Vacuum Chamber wiki.

Radhika, feel welcome to post full cooling data to this entry, or to your original - up to you!

Attachment 1: cooling_full_20210525.png
cooling_full_20210525.png
  2576   Tue May 25 13:19:41 2021 RadhikaDailyProgressCryo vacuum chamberWeekend cryo chamber cooling

[Radhika, Stephen]

Instant gratification McMaster sourcing (PO S519341, submitted earlier today)

  1. metallized PET film (off-brand Mylar) in two thicknesses
  2. PEEK sheet in a few thicknesses
  3. G10 tube with ODs matching various apertures.

Should be able to integrate some of these items soon - they arrive Thursday, and Radhika will check to see if anything works for the setup as-is, or make requests for Stephen to cut on Friday.

future Mylar source - https://www.professionalplastics.com/MYLARFILM

     - update: Radhika called Professional Plastics and they said we cannot order metallized mylar online, but we can call them back and place the order. Stephen called later and Sarah said she had to get a quote from a supplier, and will be in contact. Thickness range: 0.001-0.014 in.

future PEEK source - https://www.professionalplastics.com/PEEK_SHEET-ROD-BAR: thickness range: 0.25-2 in. (units not specified, but I assume inches).

future PEEK source - https://www.boedeker.com/Product/PEEK-Virgin-Natural: thickness range: 0.062-4 in.

uote:

exp plot tip: If you use the "grid" feature of matplotlib and plt.semilogy(), the exponentials will look like straight lines, so we can just read off the time constants with a ruler.cool


Also, as we talked about earlier today, we should make some analytical estimates for the various heat loads, and also put them into the model.

For protecting from radiation, all of the surfaces which are NOT shiny-polished should get wrapped in something shiny (UHV Al foil, with the shiny side out).

I suggest wrapping with foil:

  1. Baseplate (to keep it from radiating into the cold plate (the tapped work surface))
  2. Copper strap (after first wrapping it with a sheet of Mylar to block conduction to the foil). Al foil should be wrapped with shiny side out.
  3. Plastic spacers. wrap around the length, but not the ends.
  4. Holes in the shields where the thermal strap enters. Edge the holes with a G-10 ring/grommet to prevent touching the straps.

I attach here a photo of the radiative shielding of a Purple Pepper Plant (PPP), to reduce the radiative coupling to the environment. This prevents the soil from drying out in the sun so fast.

 

  2575   Mon May 24 16:49:17 2021 ranaDailyProgressCryo vacuum chamberWeekend cryo chamber cooling

exp plot tip: If you use the "grid" feature of matplotlib and plt.semilogy(), the exponentials will look like straight lines, so we can just read off the time constants with a ruler.cool


Also, as we talked about earlier today, we should make some analytical estimates for the various heat loads, and also put them into the model.

For protecting from radiation, all of the surfaces which are NOT shiny-polished should get wrapped in something shiny (UHV Al foil, with the shiny side out).

I suggest wrapping with foil:

  1. Baseplate (to keep it from radiating into the cold plate (the tapped work surface))
  2. Copper strap (after first wrapping it with a sheet of Mylar to block conduction to the foil). Al foil should be wrapped with shiny side out.
  3. Plastic spacers. wrap around the length, but not the ends.
  4. Holes in the shields where the thermal strap enters. Edge the holes with a G-10 ring/grommet to prevent touching the straps.

I attach here a photo of the radiative shielding of a Purple Pepper Plant (PPP), to reduce the radiative coupling to the environment. This prevents the soil from drying out in the sun so fast.

Attachment 1: PPP.jpg
PPP.jpg
  2574   Mon May 24 15:47:30 2021 RadhikaDailyProgressCryo vacuum chamberWeekend cryo chamber cooling

Today I pulled the cyro chamber cooling data from the temperature controller. Cooling started on Friday 5/21 around 1:45pm. 

The final temperatures reached at ~3:15pm today were:

outer shield: 253 K

inner shield: 168 K 

heater (off): 151 K 

workpiece: 150 K

The temperature curves (see attached) seem to be leveling off, so I'm not sure things will get too much colder. In the meantime I've reinserted the USB to resume temperature logging. 

Attachment 1: cooling.png
cooling.png
  2573   Fri May 21 11:15:12 2021 StephenDailyProgressCryo vacuum chamberTroubleshooting cryocooler - power cable not plugged in

[Radhika, Stephen remote]

After leaving the cryocooler's compressor running overnight, Radhika found all RTDs reporting room temperature. The noises coming from the compressor were normal, and all operating conditions were consistent with QIL/2504. However, the cryocooler was silent (valve motor not starting).

It turned out that the cryocooler power cable, unplugged during installation efforts (pictured below), had not been reattached. After turning off the compressor, plugging in the power cable, and turning on the compressor, the cryocooler began making normal noises and apparently operating normally. Radhika reformatted the USB drive collecting data from the temperature controller. 

Cooldown began at 1:45pm today (Friday). We will check in again on Monday.

During these troubleshooting efforts, we referred to the Cryocooler and Compressor manuals, found at the QIL Cryo Vacuum Chamber Wiki.

Attachment 1: IMG_1662.jpeg
IMG_1662.jpeg
Attachment 2: IMG_1664.jpeg
IMG_1664.jpeg
  2572   Thu May 20 16:57:32 2021 KojiElectronicsGeneralKeithley Source Meter returned to Downs

I've returned the Keithley Source Meter unit
- The unit (Keithley 2450?2460?)
- A power cable
- A pair of banana clips
- the transistor test fixture & triax cable/connectors
 

Attachment 1: P_20210520_154439.jpg
P_20210520_154439.jpg
Attachment 2: P_20210520_154505.jpg
P_20210520_154505.jpg
Attachment 3: P_20210520_154523.jpg
P_20210520_154523.jpg
  2571   Thu May 20 16:40:09 2021 RadhikaDailyProgressCryo vacuum chamberDamaged workpiece, but otherwise ready to pump down

Following up from Stephen's last post. Today I completed the outstanding tasks he outlined, with the exception of connecting a com cable to record and trend cooldown from the temperature controller. For today's cooldown we are still using the usb flash drive.

The RTD connected to the workpiece (spring clamp) took some wrestling to get stable temp readout, so I had to reclamp it. Other than that, I was able to close up the radiation shields' lids and the chamber lid straight away.

I initially tried pumping down but the backing pump was very loud, not reaching full speed, and pressure wasn't decreasing. Stephen realized the KF joint on the side of the chamber was never sealed up (just wrapped in foil) --> obvious major leak. We didn't have any blanks to seal it up, so I replaced the KF port with a blank flange at the conflat joint.

Ready to pump down once more, I ran into an error message from the turbo pump but performed a power cycle and it disappeared. Pumped down and reached a millitorr before turning on cryocooler.

I plan to pop by tomorrow morning before the cryo meeting and can share consequent updates then.

 

Quote:

[Radhika, Stephen]

Good progress toward pumping down, with a setback (impact unknown while we reach out to Karthik).

  • Made final connection of copper braid to baseplate. With the current length, we must pull the folded-over braid completely taught. We are shorting on the aperture of the inner shield, but we appear to have clearance from the vacuum skin (checked by hand at the cryocooler joint, checked by eye at the chamber) and the outer shield. We should be able to resolve these issues with the plan for a long rigid copper bar interfacing directly to the cold head, plus a short copper rope thermal linkage.
  • Completed vacuum flange connections at cryocooler and vacuum gauge on T.
    • See IMG_8743 for overview of current cryocooler connection.
  • Installed Heater on workpiece holder using 2x #4-40 screws, with an indium gasket underneath. Need to check whether we adequately compressed the indium (needs a certain pressure to flow into surface microroughness), but we didn't have any flexibility in the position or bolting arrangement - something to consider for future sample holders.
  • Connected RTDs at Inner Shield (cryo varnish), Outer Shield (cryo varnish), Heater (cryo varnish), and Workpiece (spring clamp). Heater RTD Kapton Tape separating soldered leads needed to be separated, as a short was found at those bare leads. Workpiece RTD connected directly to the workpiece holder since we want to witness the temperature of the sample as directly as possible, and the RTD mounted to the outside of the heater will be overly-sensitive to heat input, yielding a very different time constant from the workpiece (and an inaccurate witness to sample temperature).
    • See IMG_8745 for overview of internal configuration, including RTD positions and thermal linkage to baseplate.
  • Outer and Inner radiation shields were aligned to the copper braid connection to avoid shorting, then to the optical beam path by eye.
  • While making final tightening of workpiece holder, Stephen's allen key slipped, and the installed silicon cantilever was whacked and fractured. The workpiece holder will need to be registered, and the workpiece will need to be swapped out.

The following is the list of remaining actions before we have cooldown data:

  1. Replace workpiece (may require removal of workpiece holder, may require removal and reconnection of RTDs
  2. Check alignment of radiation shield apertures, confirming outer radiation shield is still standing on the 3x G10 spacers
  3. Confirm conductivity of RTDs, both by eye and by temperature controller readout.
  4. Add radiation shield lids, including bolts to inner radiation shield lid.
  5. Bolt down top chamber lid, incremental torquing 1/4-turn at a time with metal-to-metal end result (to achieve full compression of o-ring).
    1. Would be good to wipe down outer surface of o-ring using a wiper and IPA, since there have been a lot of on/off moves since we last pumped down, and we could have picked up particulate introducing a leak path. Check by eye after wipe down.
  6. Pump down system and confirm that we have no leaks (tightening new CF joints would be the first resolution, if there are any issues)
    1. Gauge cables to controller/readout are currently disconnected, need to reconnect.
    2. Can we collect and trend pumpdown data yet from the gauge controller? Might be worth setting this up, if the com cable is already in place as it appears to be.
  7. Turn on cryocooler once vacuum pressure reaches tens of milliTorr or better (should be an hour or two).
    1. Need to start USB datalogging at temperature controller.
    2. RTD names might need updating for consistency with current mounting.
    3. Can we collect and trend pumpdown data yet from the temperature controller? Might be worth setting this up, if the com cable is already in place as it appears to be.

We are both working on adding all of our photos to the photo dump at the ligo.wbridge QIL Cryostat Photo Album. We will then collaborate to add some of the most interesting images to this log!

 

Attachment 1: IMG_1646.jpeg
IMG_1646.jpeg
Attachment 2: IMG_1651.jpeg
IMG_1651.jpeg
Attachment 3: IMG_1654.jpeg
IMG_1654.jpeg
Attachment 4: IMG_1661.jpeg
IMG_1661.jpeg
  2570   Wed May 19 08:40:47 2021 StephenDailyProgressCryo vacuum chamberDamaged workpiece, but otherwise ready to pump down

[Radhika, Stephen]

Good progress toward pumping down, with a setback (impact unknown while we reach out to Karthik).

  • Made final connection of copper braid to baseplate. With the current length, we must pull the folded-over braid completely taught. We are shorting on the aperture of the inner shield, but we appear to have clearance from the vacuum skin (checked by hand at the cryocooler joint, checked by eye at the chamber) and the outer shield. We should be able to resolve these issues with the plan for a long rigid copper bar interfacing directly to the cold head, plus a short copper rope thermal linkage.
  • Completed vacuum flange connections at cryocooler and vacuum gauge on T.
    • See IMG_8743 for overview of current cryocooler connection.
  • Installed Heater on workpiece holder using 2x #4-40 screws, with an indium gasket underneath. Need to check whether we adequately compressed the indium (needs a certain pressure to flow into surface microroughness), but we didn't have any flexibility in the position or bolting arrangement - something to consider for future sample holders.
  • Connected RTDs at Inner Shield (cryo varnish), Outer Shield (cryo varnish), Heater (cryo varnish), and Workpiece (spring clamp). Heater RTD Kapton Tape separating soldered leads needed to be separated, as a short was found at those bare leads. Workpiece RTD connected directly to the workpiece holder since we want to witness the temperature of the sample as directly as possible, and the RTD mounted to the outside of the heater will be overly-sensitive to heat input, yielding a very different time constant from the workpiece (and an inaccurate witness to sample temperature).
    • See IMG_8745 for overview of internal configuration, including RTD positions and thermal linkage to baseplate.
  • Outer and Inner radiation shields were aligned to the copper braid connection to avoid shorting, then to the optical beam path by eye.
  • While making final tightening of workpiece holder, Stephen's allen key slipped, and the installed silicon cantilever was whacked and fractured. The workpiece holder will need to be registered, and the workpiece will need to be swapped out.

The following is the list of remaining actions before we have cooldown data:

  1. Replace workpiece (may require removal of workpiece holder, may require removal and reconnection of RTDs
  2. Check alignment of radiation shield apertures, confirming outer radiation shield is still standing on the 3x G10 spacers
  3. Confirm conductivity of RTDs, both by eye and by temperature controller readout.
  4. Add radiation shield lids, including bolts to inner radiation shield lid.
  5. Bolt down top chamber lid, incremental torquing 1/4-turn at a time with metal-to-metal end result (to achieve full compression of o-ring).
    1. Would be good to wipe down outer surface of o-ring using a wiper and IPA, since there have been a lot of on/off moves since we last pumped down, and we could have picked up particulate introducing a leak path. Check by eye after wipe down.
  6. Pump down system and confirm that we have no leaks (tightening new CF joints would be the first resolution, if there are any issues)
    1. Gauge cables to controller/readout are currently disconnected, need to reconnect.
    2. Can we collect and trend pumpdown data yet from the gauge controller? Might be worth setting this up, if the com cable is already in place as it appears to be.
  7. Turn on cryocooler once vacuum pressure reaches tens of milliTorr or better (should be an hour or two).
    1. Need to start USB datalogging at temperature controller.
    2. RTD names might need updating for consistency with current mounting.
    3. Can we collect and trend pumpdown data yet from the temperature controller? Might be worth setting this up, if the com cable is already in place as it appears to be.

We are both working on adding all of our photos to the photo dump at the ligo.wbridge QIL Cryostat Photo Album. We will then collaborate to add some of the most interesting images to this log!

Attachment 1: IMG_8743.JPG
IMG_8743.JPG
Attachment 2: IMG_8745.JPG
IMG_8745.JPG
  2569   Mon May 17 09:52:37 2021 RadhikaDailyProgressCryo vacuum chamberLarge cyro chamber progress

[Stephen, Radhika]

As of last Thursday (5/13), the new vacuum tubes had been selected and bolted in. We fixed the copper braid to the cryo-cooler: the braid was folded in half, with the loose ends bolted to the cryo-cooler and the folded end fed through the vacuum tubes into the chamber. The folded loop was bolted down to the baseplate. The copper braid was pulled tight and thus has no slack. Aluminum sheets were used to wrap frayed areas of the copper braid to prevent shorting to the tube walls, though this needs to be revisited. We also still need to address shorting to the inner/outer radiation shields.

Attachment 1: IMG_1615.jpeg
IMG_1615.jpeg
Attachment 2: IMG_1625.jpeg
IMG_1625.jpeg
Attachment 3: IMG_1634.jpeg
IMG_1634.jpeg
Attachment 4: IMG_1641.jpeg
IMG_1641.jpeg
  2568   Wed May 12 15:43:30 2021 AidanSummary2um PhotodiodesChamber is leaking

I tried pumping down the JPL PD chamber to test the new PD at cryo temperatures. Unfortunately, the chamber can;t get past about 6E-3 Torr with the pump on. As soon as I turned off the pump the pressure rose to around 2 Torr over 20 minutes or so.

I extricated the chamber from the pedestals, flipped it and removed the bottom plate. I cleaned the O-ring with isopropanol and wiped down the mating surface on the chamber (also with iso). I replaced the plate and tightened the screws. Then I returned the chamber to the table and reconnected it to the vacuum system. I tried pumping down once again but I saw pretty much exactly the same situation as before (pressure bottoming out around 6E-3 Torr and then rising quickly again when the pump was turned off).

I guess it's possible that the O-ring is damaged - although I couldn't see anything obivous. We didn't mess around with the viewport (when we replaced the diode a few weeks ago) so I'm hoping there is no issue there.

  2567   Tue May 11 19:23:51 2021 RadhikaDailyProgress2um PhotodiodesADC noise level spectrum

It is important to characterize the noise levels of all instruments used in the current PD testing setup. We generally expect ~5uV/rHz of ADC input noise. Verifying/correcting this value will be key to ensuring that our overall gain is enough to amplify various signals above the ADC noise floor.

I terminated the input to ADC channel 31 with a 50-ohm BNC terminator. I used diaggui to generate the resulting amplitude spectra, with 0.03 BW (attached). To convert counts to volts, I took a range of 20V divided by 2^16 counts, resulting in a scaling of 3e-4 V/count. I plan to conduct another test to confirm this value (feeding a known voltage and comparing to the output). In the meanwhile, the resulting noise level consistent with our expectation of a few uV/rHz.

Attachment 1: adcnoise_210511.xml.gz
Attachment 2: adcnoise.png
adcnoise.png
  2566   Mon May 10 15:38:36 2021 ranaSummary2um PhotodiodesKeithley connections

Note that the back panel connectors are Triax, not the usual Coax.

  2565   Sun May 9 11:14:30 2021 PacoDailyProgressCryo vacuum chamberCooling power of current setup

Going back to a 1D heat transfer model, and matching Stephen's numbers for the area and length of the thermal strap, I confirm that the conductive power of a single strap is indeed heavily constraining the cryocooler capacity. For my simulation the peak power for a 0.5m copper strap with area 6.71e-5 m^2 is 1.06 Watts with an average of a few hundred mW during the cooldown.

The main difference with respect to Stephen's numbers is that I account for temperature dependent conductivity and heat capacity, include a radiative sink (surrounding vacuum tube at room temp), and take the strap to be made of RRR 500 Cu.

Attached is the predicted temperature at the end of the strap as a function of time when the operating point of the cryocooler is set to 123 K. Note the cooldown delay caused by the single half-meter strap (with respect to the cryocooler cold head).

Attachment 1: strap_1D_heat_transfer.pdf
strap_1D_heat_transfer.pdf
  2564   Wed May 5 00:34:14 2021 ranaSummary2um PhotodiodesUpdated PD testing schematic / measurement table

Looks very clear, thanks. I guess the next thing to do is

  1. ask if this will work for all the various PDs we want to test,
  2. is it good enough for all our requirements, and then we
  3. draw a new diagram for the new setup, incorporating what to keep and what circuit to make ourselves
  2563   Mon May 3 17:32:29 2021 RadhikaSummary2um PhotodiodesUpdated PD testing schematic / measurement table

Attached:

- Updated schematic of the current PD testing setup, including noise levels for current electronics

- Table of desired measurements for new setup, with expected signal levels, accuracy, and readout values

Attachment 1: QIL_PD_testing.graffle
Attachment 2: QIL_PD_testing.pdf
QIL_PD_testing.pdf
Attachment 3: Screen_Shot_2021-05-03_at_17.16.20.png
Screen_Shot_2021-05-03_at_17.16.20.png
  2562   Mon May 3 16:56:59 2021 StephenDailyProgressCryo vacuum chamberCooling power of current setup

StephenA

This log investigates cooling through our current planned copper braid connection (which is standing in for an intended rigid bar linkage that is WIP)

The question is, can we get [cooling power of cryocooler] out of our baseplate through this copper braid?

Copper Braid

Cooner Wire P/N NER 7710836 BOF (oxygen free copper)

  • AWG Size 2/0
  • Circular Mil Area 132300 (conversion: .104 in^2 = 6.71e-5 m^2) - note that a 1 cm x 1 cm bar would have an area of 1e-4 m^2
  • No. of Wires 5292
  • Wire AWG Size 36
  • Construction 7x7x108/36
  • Nominal Diameter .483"
  • Pounds Per MFT 433.

ref. https://www.coonerwire.com/flexible-wire-rope/

Cryocooler

Sumitomo CH-104 (manual from Wiki) has 77K coldhead cooling capacity of 34 W, and from the quote, 50K cooling capacity of just under 40 W.

Adequate cooling power of this setup depends on the radiative heat load and conductive losses; for our purposes, we can imagine that tens of Watts will be needed, and circle back to more precise heat budgeting.

Conductive Heat Transfer

Q = A / L * (Uint_T2 - Uint_T1)

Uint_T = the integral of thermal conductivity between T and 4K, see below table [ETP OFE Copper, W/m]. Note these are values from literature not from our copper braid's spec sheet (no such properties available from vendor).

Table of Thermal Conductivity integral values, between T and 4K. Unit = W/m. Source: Ekin, Appendix 2.1

               20K = 14000, 40K = 40600, 50K  = 50800, 60K = 58700, 70K = 65100, 80K = 70700, 100K = 80200, 120K = 89100, 140K = 97600

A = 6.71e-5 m^2

L = 0.5 m (estimate)

T2 = 123 K (intended workpiece temperature)

T1 = ? (coldhead temperature, unknown, we will pick a value and calculate)

Q(T1_80K) = 6.71e-5 m^2 / 0.5 m * (89100 W/m - 70700 W/m) = 2.46 W

Q(T1_20K) = 6.71e-5 m^2 / 0.5 m * (89100 W/m - 14000 W/m) = 10.07 W

Conclusion

It appears that the copper braid's capacity for conductive heat transfer will constrain the tens of Watts of cryocooler capacity. This is even before we consider imperfections in the clamping interfaces and similar real losses.

Fixes for this constraint might involve adding parallel linkages (increasing area) or shortening the strap length.

It would be interesting to compare this to the anticipated capacity of the flexible strap in the original design - future work.
 

Attachment 1: IMG_8665.JPG
IMG_8665.JPG
  2561   Fri Apr 30 08:09:08 2021 StephenDailyProgressCryo vacuum chamberSimplifying cryocooler connections is WIP

RadhikaB, StephenA

Two sessions this week were spent working toward simplification of the cryocooler connection.

We needed to order a couple of off the shelf vacuum fittings to complete the intended design - image attached.

Still WIP!

Attachment 1: D2100320_cryostat_update_new_design_20210430.png
D2100320_cryostat_update_new_design_20210430.png
  2560   Wed Apr 21 16:45:01 2021 AidanLab InfrastructureCamerasTelephoto and macro lenses for smartphones

I purchased a set of telephoto and macro lenses for the lab. They're stored in the tool cabinet.

 

Attachment 1: IMG_2194.jpg
IMG_2194.jpg
Attachment 2: IMG_2195.jpg
IMG_2195.jpg
Attachment 3: IMG_2196.jpg
IMG_2196.jpg
  2559   Wed Apr 21 09:42:35 2021 Stephen, ranaThings to BuyCryo vacuum chamberReimagining QIL Cryo Vacuum Chamber

WIP log entry - working on getting all of our ideas down on the page, then will sort and elaborate.


We met to discuss a range of topics relating to the path ahead for the Cryo Vacuum Chamber. This reconsideration of the current state of things is necessary as the chamber needs to become the workhorse for PD characterization efforts soon, in addition to a range of other tests (large suspension tests will be conducted in a different chamber, yet to be designed)

  1. Pumping station should be moved away from table, with long roughing lines perhaps coming down from above using some ceiling-mounted cable tray or similar.
  2. Primary pumping line to chamber is large and overkill in terms of conductance. Can move to a longer, more flexible, smaller diameter connection (may need to adapt using CF zero-length flange).
  3. Pumping of the external volume may be managed by valve arrangement and direct connection of both volumes to the turbo pump. Valve out the external volume once the pressure is low, and the pressure should hold well enough for conductive losses to be minimal.
  4. Viewports seem suitable, no issues throughout. I learned about common coating behavior, namely that reflectance is generally at half the wavelength of transmission, so if I see a green reflection it suggests transmission at IR. Neat!
  5. Cabling into chamber for temp sensors is pretty scary. Noticed the kink in the cable bundle caused by the flexible part of the cable extending longer than the grip of the connector's stiffening. Needs to be reassembled with the stiff cabling under the grip (could extend the grip, shorten the flexible leads, etc.) to avoid the kink.
  6. External volume feels like a misdirected design.
    1. For general case, consider mounting the cryocooler directly to chamber. Avoid losses related to thermal linkages from cold head of cryocooler to baseplate (diagram will be supplied, Rana was especially concerned that the V clamping arrangement didn't have adequate contact area in the line contacts)
    2. For vibration-sensitive experiments, would be good to have a flexible bellows reducing vibrational energy through the vacuum skin, and flexible strap reducing vibrational energy to the baseplate.
    3.  For contamination-sensitive experiments, would be necessary to implement a feedthrough as currently, but seems overkill for current slate of experiments.
    4. The intention seems to have been to devise a scaleable solution that would work for Mariner, but we are currently very far from realizing that (cryocooler needs to be Stirling cycle for vibration, no validation yet of adequate thermal conductivity through the external volume to the baseplate, etc.)
  7. Thermal straps with Mylar shielding is not the optimal implementation. Something like a rigid copper bar provides better conductivity, and can be shielded by G10 tubing with Aluminum metallization on the OD. If a flexible connection is required (for example, vibration isolation or positional uncertainty) a thermal strap may also be shielded in this sort of conduit.
  8. Need to replace the V-groove copper connections with something with much more surface area. V-grooves are nearly 1D contact lines, so they are probably the mian cooling rate limiter at the moment. Need to get some new parts fabbed ASAP to continue working on this cryostat.
  9. Vent valve should be a leak valve with a controlled, small conductance, perhaps backed by a filter. Want to allow slow, controlled venting.
  10. Skyhook Crane should be on its cart wheels for easy relocation away from the experiments.
  11. Skyhook should be replaced by a simple hoist mounted to ceiling of enclosure. Would require stiffening of some members of the enclosure's ceiling, but would permit easier access with fewer traffic jams.
  12. Yellow solvent cabinet should be removed. Solvents should be stored under fume hood.
  13. Would be great to get a stand (ie wire shelving, but heavy duty) which could hold the compressor, hold the pumping station, and provide a single location for any other items that need to be interfaced. All connections would be routed over the walkway via a run of cable rack.
  14. Moving the chamber to the center of the table width would be helpful to opening up access to more ports. Currently located in a corner, such that only half of ports are accessible. This is an extra reason that the Sky Hook should be remounted to its wheeled base.
  15. Should put all controllers on a rack, rather than consuming optical table space. All necessary serial comm cabling could then run to this singular location. We can use the rack next to the sink which has the NIM racks. The rack is completely unused right now. Need to get some rack parts to put some shelving in there.
  16. Documentation of ongoing thoughts, design efforts, modifications, etc. can be contained at the wiki!
  17. What is currently installed? Some insights from the wiki (ie gauges, pumps, viewports) should be elevated into a comprehensive diagram with a bill of materials or similar.
  18. In the diagram need to note all instruments so that Radhika can include it in her work to interface with the DAQ. i.e. no more photos and screenshots to record data.
  19. Where am I? A floorplan for this experiment (current and planned) would be worth some time, now that we are considering specific improvements.

Will sort the above into some sort of timeline (such as short term / long term).

Ruminations about the future chamber for suspension work:

  1. Stirling cryocooler for vibration isolation
  2. Straight-sided construction (i.e. rectangular prism) for more usable footprint inside
  3. If vertical (i.e. lid) use counterbalance and hinge for easy opening without hoist or crane; might need side ports large enough for hand access, since the height will be prohibitive to reach all the way down to the baseplate.
  4. Horizontal (i.e. door) might be preferred, especially with the volume in the form of a rectangular prism. This would allow access throughout the height of the chamber.
  5. Usual vacuum vendors should be able to help bring the design from a sketch to a quote, so start those conversations.
  2558   Tue Apr 20 09:43:49 2021 RadhikaSummary2um PhotodiodesLink to PD testing master doc

Posting link to PD testing google doc here: 

https://docs.google.com/document/d/1RrtX5nqNeEOazNvT2rPxHnSqumgvYfaBrYTMMGZ-ecc/edit

  2557   Fri Apr 16 13:05:05 2021 AidanUpdate2um Photodiodes2um DC photodiode voltage weird drift with no incident power

We put the preamp output directly into a multimeter and observed the same fluctuating behavior as the DAC channel was changed.

We're bypassing the relay to see if that makes any difference. The old relay wiring (to be bypassed) is shown in the attached diagram. That didn't do anything.

We're looking at filtering the DC output by 5kHz to see if there are any resonances at higher frequencies that might go away. Changing SR560 output for AC path to DC and setting gain to 1 on that unit. Also changing gain in FM31 filter bank from 1E-3 to 1. The results are shown in the attached time series. The channels FM30 and FM31 see the same thing. The only difference is that FM31 goes through an SR560 with a 0.03Hz pole (6dB).

Success by bypassing the DAC bias voltage. We switched to a 300mV bias voltage from a function generator. Doing that removed the causal PD voltage drift induced by changing the laser diode current set voltage (see the last time series). So the issue is some weird coupling into the DAC bias voltage.

Quote:

[Aidan, Radhika, Nina]

We noticed that the DC channel readout (FM30) of the JPL A1 photodiode is drifting around. What we observe with no light on the photodiode, is the DC output drifiting around. It gets particularly bad when we apply voltage to other DAC channels.

For example, the attached plot shows the DC voltage from the photodiode as I change the set voltage to the laser diode driver. To be absolutely clear, the laser driver itself was completely powered off. I'm just varying the voltage going into the set point BNC connector on the back of it.

For reference, the set up is:

DAC (300mV bias) > relay > PD > relay  > FEMTO preamp (1000x gain) > ADC channel FM30

 

 

Attachment 1: IMG_2132.jpg
IMG_2132.jpg
Attachment 2: FM30_31_w_latter_filtered.png
FM30_31_w_latter_filtered.png
Attachment 3: IMG_2134.jpg
IMG_2134.jpg
Attachment 4: bias_bypassed.png
bias_bypassed.png
  2556   Fri Apr 16 11:27:01 2021 AidanUpdate2um Photodiodes2um DC photodiode voltage weird drift with no incident power

[Aidan, Radhika, Nina]

We noticed that the DC channel readout (FM30) of the JPL A1 photodiode is drifting around. What we observe with no light on the photodiode, is the DC output drifiting around. It gets particularly bad when we apply voltage to other DAC channels.

For example, the attached plot shows the DC voltage from the photodiode as I change the set voltage to the laser diode driver. To be absolutely clear, the laser driver itself was completely powered off. I'm just varying the voltage going into the set point BNC connector on the back of it.

For reference, the set up is:

DAC (300mV bias) > relay > PD > relay  > FEMTO preamp (1000x gain) > ADC channel FM30

 

Attachment 1: DC_2um_voltage_weirdness.png
DC_2um_voltage_weirdness.png
  2555   Wed Apr 14 09:49:38 2021 StephenDailyProgressCryo vacuum chamber 

2021.04.14 StephenA

QIL Cryo vacuum chamber cooldown was not as successful under the new configuration (radiation shielded by cylindrical outer + inner shields, cold finger thermally strapped to baseplate).

--> Karthik's Si cantilever workpiece was stable at 240 K.

--> Cold Finger was stable at 200 K - there is significant thermal loss between the cold finger and the workpiece.

--> Inner shield was stable at 250 K - seems to be somewhat decoupled from the baseplate; not very satisfied with the current state of the shielding.

Will need to re-examine some of the connections, which were not optimal (especially the improvised dog clamped strap-baseplate interface). Fabricating an adapter piece for the thermal strap which will be bolted 4x on a 2" x 2" grid. Might also look into a new thermal strap which could interface with baseplate directly.

Also will need to consider options to decouple outer shield from inner, and double check that shield orientation has no other solution (hoping there's an answer to the question, why would outer shield be coupled to baseplate?)

 

Data - cooldown 20210408 (CSV = raw, XLSX = Stephen's plots) in Box Folder [Voyager\MarinerBox\CryoEngineering\CSVlogs]

Description - 6 day cooldown. Layout described in QIL/2552. The radiation shields were installed and thermal strap was connected to baseplate. The cryocooler was turned on/off at the start/end of the data collection, and the in-vac heater was not powered on at all.

Images -

  1. IMG_8570 = starting conditions;
  2. IMG_8585 = final conditions after 144 hours;

Plots -

  1. cooldown_20210408_first_si_workpiece_with_shields_and_straps

Attachment 1: IMG_8570.JPG
IMG_8570.JPG
Attachment 2: IMG_8585.JPG
IMG_8585.JPG
Attachment 3: cooldown_20210408_first_si_workpiece_with_shields_and_straps.png
cooldown_20210408_first_si_workpiece_with_shields_and_straps.png
  2554   Tue Apr 13 11:48:04 2021 RadhikaDailyProgress2um PhotodiodesOpened cryo chamber and removed old PD

I posted a video tutorial of the diode replacement.

https://dcc.ligo.org/G2100807

Quote:

04/01/21

Aidan and I removed the old PD from the cryo chamber in order to start testing C3 (plan for tomorrow, 04/02).

Steps:

- Brought chamber up to room pressure, disconnected readout wires and vacuum pump.

- Picked up chamber and placed it upside down on makeshift support stand (see pics).

- Unscrewed mounting plate and 2 inner insulation plates to reveal mounted PD.

- Had trouble unscrewing PD mount, since the screws were very close to the PD and we had to be careful not to slip and cause damage. Started with 2 side screws, then bottom (hardest), then top.

- Successfully removed PD and put away. Placed chamber components back in place without bolting in.

- Plan is to mount PD C3 in chamber tomorrow and begin testing.

 

 

  2553   Fri Apr 9 16:23:22 2021 ranaDailyProgress2um PhotodiodesOpened cryo chamber and removed old PD

To aid in taking photos of these diodes, I put a USB microscope on Anchal's desk - you can grab it from there. I use it with mac Photo Booth, but it should be easy to use with any camera application.

Also, I recommend buying a macro lens(es) for cell phones from Amazon or B&H.  Label them with the QIL lab sticker so they don't disappear.

  2552   Fri Apr 9 09:37:26 2021 StephenDailyProgressCryo vacuum chamberStarted cooldown with shielding, workpiece Si cantilever

2021.04.08, StephenA

Karthik had completed in-chamber alignment efforts during a prior visit. In air alignment also completed following viewport move.

0) Removed lid for access to chamber.

--> posted demo video to ligo.wbridge QIL Cryostat HowTo Playlist.

1) Mounted RTDs to final positions - locations are Heater (cryo varnish+cigarette paper, pictured in IMG_8558 curing under weight of upsidedown bolt), Inner Shield (cryo varnish+cigarette paper, pictured in IMG_8559), Cold Finger (spring clamp), and Workpiece (spring clamp).

--> Final chamber layout may be viewed in IMG_8562

--> Note that Karthik's Si cantilever, mounted vertically in the right of the image, is NOT bolted down to the baseplate (just located on baseplate by dog clamps, held down via gravity). This will need to be investigated to enable workpiece cooling.

2) Installed radiation shield lids - no bolts to expedite the process and to see if there is any bulk motion during pumpdown and thermal cycling.

--> note that the lid for the outer radiation shield seems to interface with the current shield orientation perfectly; if there was a mismatch, it would point toward the inverted orientation being intended, but this seemed pretty definitive.

3) Installed the cryostat lid - final positioning and alignment made easier by teflon rails!

--> posted demo video to ligo.wbridge QIL Cryostat HowTo Playlist.

4) Pumped down - single button press to turn on pumping station.

--> note that it took about 1 hour for both gauges to reach a few mTorr.

5) Confirmed function of heater - set PID setpoint to 350 K and enabled outputs, observed temperature rise in heater RTD.

--> note that PID autotuning should be done at steady state with workpiece RTD, before enabling outputs again!

6) Turned on cryocooler - flip power lever and turn on green system switch.

--> start time was 10 am.

7) Started temperature datalogging to USB - press dull red indicator dot on upper right corner of CTC-100 once, and note that indicator is now bright red.

8) Remaining photos posted to the ligo.wbridge QIL Cryostat Photo Album

Attachment 1: IMG_8558.JPG
IMG_8558.JPG
Attachment 2: IMG_8559.JPG
IMG_8559.JPG
Attachment 3: IMG_8562.JPG
IMG_8562.JPG
  2551   Mon Apr 5 18:50:54 2021 RadhikaSummary Current PD testing schematic

I'm attaching my rough first draft of the QIL photodiode testing schematic. Please provide comments for fixes/improvement!

Attachment 1: QIL_PD_testing.jpg
QIL_PD_testing.jpg
Attachment 2: QIL_PD_testing.graffle
  2550   Thu Apr 1 20:54:57 2021 RadhikaDailyProgress2um PhotodiodesOpened cryo chamber and removed old PD

04/01/21

Aidan and I removed the old PD from the cryo chamber in order to start testing C3 (plan for tomorrow, 04/02).

Steps:

- Brought chamber up to room pressure, disconnected readout wires and vacuum pump.

- Picked up chamber and placed it upside down on makeshift support stand (see pics).

- Unscrewed mounting plate and 2 inner insulation plates to reveal mounted PD.

- Had trouble unscrewing PD mount, since the screws were very close to the PD and we had to be careful not to slip and cause damage. Started with 2 side screws, then bottom (hardest), then top.

- Successfully removed PD and put away. Placed chamber components back in place without bolting in.

- Plan is to mount PD C3 in chamber tomorrow and begin testing.

 

Attachment 1: IMG_1834.jpg
IMG_1834.jpg
Attachment 2: IMG_1837.jpg
IMG_1837.jpg
Attachment 3: IMG_1838.jpg
IMG_1838.jpg
Attachment 4: IMG_1842.jpg
IMG_1842.jpg
Attachment 5: IMG_1841.jpg
IMG_1841.jpg
Attachment 6: IMG_1844.jpg
IMG_1844.jpg
Attachment 7: IMG_1845.jpg
IMG_1845.jpg
  2549   Thu Apr 1 20:33:42 2021 StephenDailyProgressCryo vacuum chamberInstallation of Radiation Shields

2021.03.17, StephenA

1. Radiation Shields located (in TCS lab), unwrapped, fitted up.

Location in TCS Lab - IMG_8351

Removed lid and placed adjacent to chamber (cleared a little space, used 3 plastic flange covers to make nonmarring surface safe for lid and o-ring - IMG_8353

Fit up as installed - IMG_8354

Comments on the fit up - I looked at all of the apparent sources for insights into Rahul's original design intent - QIL elog 2276, DCC T1800308-v1, wiki for Cryo Vacuum Chamber. It appears that Rahul never decoupled the upper outer radiation shield from the cold plate, which seems like a strange omission. Chris and Raymond also appear to have been wrapping their heads around the intended layout, they came up with the fit up in QIL elog 2429 and sketch from QIL elog 2430. I will revisit their sketch at a future opportunity, but I went with something closer to 2429 as I was concerned about the height misalignments they described. Note that the height misalignment appears in Rahul's T1800308 CAD (see T1800308-v1 screenshot) so who knows what's "correct". I'll work on finalizing D2100310 CAD with radiation shield to capture the true current dimensions and fit up, to hopefully avoid such issues in the future.

2. Radiation Shields installed in Cryostat. Sequence was important here, as were a couple of improvised solutions to shortcomings of the existing parts.

Dog Clamps placed on bottom plate (to stand off bottom radiation shield bottom lid; not pictured, I think I placed some alumina washers on the dog clamps as well, not sure though anymore!). Also pictured are the usual PEEK legs for cold plate - IMG_8355

Bottom radiation shield bottom lid placed on dog clamps spacer, and bottom radiation shield cylinder placed on bottom lid - IMG_8356. Seems likely that the bottom radiation shield would be better configured upside-down.

Bolted cold plate down onto legs, with cold plate decoupled from bottom radiation shield - IMG_8357

Outer radiation shield installed and inner radiation shield installed (both needed to be tipped into place gingerly, but both cleared the cold finger cylinder with the flange removed. The heater also passed through the apertures successfully - IMG_8358

2x Alumina washers placed under outer radiation shield, inner radiation shield on cold plate - IMG_8374

Cold Flange reinstalled, though one of the brass SHCS was sheared - this was due to over torque, with 20 in*lb applied by mistake. Correct torque is 10 in*lb. The remaining 3 bolts were tightened to 10 in*lb. - IMG_8360

Top view of radiation shield apertures and cold plate grid - IMG_8375

Thermal strap interface to cold plate - dog clamps required due to strange spacing of clearance holes. - IMG_8376

 

 

Attachment 1: IMG_8351.JPG
IMG_8351.JPG
Attachment 2: IMG_8353.JPG
IMG_8353.JPG
Attachment 3: IMG_8354.JPG
IMG_8354.JPG
Attachment 4: t1800308-v1_cryostat_step_section_screenshot.png
t1800308-v1_cryostat_step_section_screenshot.png
Attachment 5: IMG_8355.JPG
IMG_8355.JPG
Attachment 6: IMG_8356.JPG
IMG_8356.JPG
Attachment 7: IMG_8357.JPG
IMG_8357.JPG
Attachment 8: IMG_8358.JPG
IMG_8358.JPG
Attachment 9: IMG_8374.JPG
IMG_8374.JPG
Attachment 10: IMG_8360.JPG
IMG_8360.JPG
Attachment 11: IMG_8375.JPG
IMG_8375.JPG
Attachment 12: IMG_8376.JPG
IMG_8376.JPG
  Draft   Thu Apr 1 20:28:26 2021 RadhikaDailyProgress2um PhotodiodesOpened cryo chamber and removed old PD

04/01/21

Aidan and I removed the old PD from the cryo chamber in order to start testing C3 (plan for tomorrow, 04/02).

Steps:

- Brought chamber up to room pressure, disconnected readout wires and vacuum pump.

- Picked up chamber and placed it upside down on makeshift support stand (see pics).

- Unscrewed outer and inner insulation plates.

 

  2547   Thu Apr 1 18:42:55 2021 Stephen DailyProgressCryo vacuum chamberPreparations for Q measurements

2021.03.30, StephenA

1. Viewport swap to nozzle that is not occluded by cryo shield = complete. All bolts on both Active Ion Gauge and Viewport have been torqued gradually (about a half turn at a time, around the clock dial) until the conflat seal was metal-to-metal. Periscope on damped optical rod was rotated to make room for replacement.

Before viewport swap - IMG_8487

After viewport swap - IMG_8502

2. Cryo RTD repair = complete. Two RTDs had been damaged during prior mounting efforts by me. I was able to repair the clamped RTD at the single damaged solder joint. I was able to repair the former Al-Block RTD by replacing the RTD element, and making a new direct attachment (not preloaded, not varnished to the aluminum block anymore)

Materials and set-up for solder repair - IMG_8491

Repaired Clamp-2 RTD - IMG_8494

Damaged Al-Block RTD - IMG_8492 (note short length between kapton strain relief and aluminum block was not ideal, one lead had already fractured and the second soon followed at the slightest touch)

Repaired, remounted Al-Block RTD - IMG_8495 (heater sandwiched underneath threaded adapter, clamp threaded into adapter, sandwiching RTD at top plane)

Remounted Clamp-2 RTD - IMG_8496 (RTD clamped at cold flange, strap is mounted)

Remaining Clamp-1 and Varnish RTDs are free - IMG_8497

Current readouts of CTC-100 controller, with repaired RTDs now behaving (note need to rename the Al-Block RTD) - IMG_8501

3. Next steps:

 - Karthik to install clamps, align in-air relay, and confirm positition of radiation shield aperture.

 - Remaining free RTDs to be mounted; current RTDs are mounted at Heater and Cold Flange, would be good to mount RTD at Work Piece/Clamp and at Outer Radiation Shield.

 - Radiation shield lids to be installed (might be easiest to install Outer Radiation Shield RTD after installing lid)

 - Mount lid, install bolts, pump down, turn on cryo cooler, the usual!

Attachment 1: IMG_8487.JPG
IMG_8487.JPG
Attachment 2: IMG_8502.JPG
IMG_8502.JPG
Attachment 3: IMG_8491.JPG
IMG_8491.JPG
Attachment 4: IMG_8494.JPG
IMG_8494.JPG
Attachment 5: IMG_8492.JPG
IMG_8492.JPG
Attachment 6: IMG_8495.JPG
IMG_8495.JPG
Attachment 7: IMG_8496.JPG
IMG_8496.JPG
Attachment 8: IMG_8497.JPG
IMG_8497.JPG
Attachment 9: IMG_8501.JPG
IMG_8501.JPG
ELOG V3.1.3-